Gabrielle is 15 years younger than Mikhail. The sum of their ages is 81 . What is Mikhail's age?
Answer by nerdybill(7384)

Gabrielle is 15 years younger than Mikhail. The sum of their ages is 81 . What is Mikhail's age?
.
Let x = Mikhail's age
then
x-15 = Gabrielle's age
.
x + x-15 = 81
2x-15 = 81
2x = 96
x = 48 years old